OpenStudy (anonymous):

Find f(-2) if f(x) + -3x - 1

OpenStudy (anonymous):

You need to put the value in the function, ie replace x with -2

OpenStudy (anonymous):

f(-2)= -3(-2) -1 f(-2)=6-1 f(-2)=5

OpenStudy (anonymous):

Thanks !

OpenStudy (anonymous):

you're most welcome ^_^